# Deep Modularity Networks with Diversity–Preserving Regularization

### v1

@@ B.5 Coauthor Physics
993 DMoN 12.61 60.14 44.16 46.45
550 DMoN 14.36 63.58 51.07 45.68
243 DMoN 13.84 63.82 56.68 49.40
16 DMoN 13.81 63.78 57.00 49.64
716 DMoN 13.91 63.70 55.59 49.16
383 DMoN 13.84 63.08 48.79 45.27
277 DMoN 14.06 63.70 46.72 39.63
274 DMoN 13.84 63.93 53.70 46.41
188 DMoN 13.94 63.70 55.66 49.31
796 DMoN 13.89 64.04 54.63 47.03
Mean 13.81 63.35 52.40 46.80
Variance 0.00 0.01 0.20 0.09
@@
993 DMoN–DPR (D) 16.08 60.64 49.75 44.05
550 DMoN–DPR (D) 15.42 56.02 50.59 52.21
243 DMoN–DPR (D) 15.78 58.21 58.58 53.21
16 DMoN–DPR (D) 15.91 54.69 53.21 54.59
716 DMoN–DPR (D) 15.85 57.35 50.14 49.28
383 DMoN–DPR (D) 18.82 55.05 51.96 48.46
277 DMoN–DPR (D) 19.89 54.78 50.10 47.08
274 DMoN–DPR (D) 15.03 56.06 55.73 54.53
188 DMoN–DPR (D) 13.71 59.75 61.71 54.99
796 DMoN–DPR (D) 18.02 56.74 58.17 51.22
Mean 16.45 56.93 53.99 50.96
Variance 0.04 0.04 0.18 0.13
@@
993 DMoN–DPR (V) 12.00 58.02 45.50 45.54
550 DMoN–DPR (V) 12.44 57.44 53.99 54.11
243 DMoN–DPR (V) 13.82 61.60 59.00 51.98
16 DMoN–DPR (V) 13.64 58.69 58.90 56.75
716 DMoN–DPR (V) 14.06 61.38 56.03 51.89
383 DMoN–DPR (V) 14.42 63.20 52.72 45.15
277 DMoN–DPR (V) 12.96 57.38 52.20 53.30
274 DMoN–DPR (V) 12.88 57.23 54.09 56.20
188 DMoN–DPR (V) 13.95 57.14 52.27 54.06
796 DMoN–DPR (V) 14.29 61.71 55.55 49.94
Mean 13.45 59.38 54.03 51.89
Variance 0.01 0.05 0.15 0.16
@@
993 DMoN–DPR (E) 12.88 58.41 44.16 45.67
550 DMoN–DPR (E) 12.07 57.62 54.04 53.83
243 DMoN–DPR (E) 13.79 62.02 58.66 51.38
16 DMoN–DPR (E) 13.77 60.22 58.66 53.25
716 DMoN–DPR (E) 14.41 61.97 56.14 50.58
383 DMoN–DPR (E) 13.88 58.47 53.74 53.85
277 DMoN–DPR (E) 12.85 57.66 52.26 52.98
274 DMoN–DPR (E) 15.10 59.77 51.39 45.72
188 DMoN–DPR (E) 14.12 62.20 56.74 50.19
796 DMoN–DPR (E) 15.15 62.10 45.87 43.09
Mean 13.80 60.04 53.17 50.05
Variance 0.01 0.04 0.25 0.15
@@
993 DMoN–DPR (DV) 14.08 60.60 47.53 46.00
550 DMoN–DPR (DV) 14.07 55.94 54.18 57.79
243 DMoN–DPR (DV) 13.03 58.15 60.48 59.90
16 DMoN–DPR (DV) 12.03 55.51 54.29 57.77
716 DMoN–DPR (DV) 12.54 56.50 52.40 57.28
383 DMoN–DPR (DV) 13.82 56.10 52.59 57.34
277 DMoN–DPR (DV) 13.96 56.84 53.28 57.82
274 DMoN–DPR (DV) 12.40 56.41 52.47 57.72
188 DMoN–DPR (DV) 14.36 56.80 60.08 57.68
796 DMoN–DPR (DV) 14.34 55.61 53.64 54.26
Mean 13.46 56.85 54.09 56.36
Variance 0.01 0.02 0.14 0.15
@@
993 DMoN–DPR (DE) 15.49 55.62 48.15 53.66
550 DMoN–DPR (DE) 15.31 55.21 52.10 56.57
243 DMoN–DPR (DE) 13.74 56.71 59.84 58.07
16 DMoN–DPR (DE) 12.99 55.20 53.27 56.69
716 DMoN–DPR (DE) 14.00 56.38 59.34 58.29
383 DMoN–DPR (DE) 13.94 56.83 55.09 56.84
277 DMoN–DPR (DE) 14.97 54.26 54.14 60.85
274 DMoN–DPR (DE) 12.57 55.85 57.05 59.56
188 DMoN–DPR (DE) 12.99 56.54 58.60 58.86
796 DMoN–DPR (DE) 16.16 58.41 49.49 49.99
@@
993 DMoN–DPR (VE) 8.50 36.32 14.32 44.25
550 DMoN–DPR (VE) 13.12 57.28 52.98 56.87
243 DMoN–DPR (VE) 14.64 59.71 55.12 51.50
16 DMoN–DPR (VE) 13.92 58.45 62.37 58.80
716 DMoN–DPR (VE) 13.75 57.27 52.06 51.47
383 DMoN–DPR (VE) 14.04 57.77 56.72 58.23
277 DMoN–DPR (VE) 13.04 57.90 54.22 57.55
274 DMoN–DPR (VE) 12.10 56.56 51.63 56.60
188 DMoN–DPR (VE) 10.40 55.47 55.62 62.42
796 DMoN–DPR (VE) 15.01 60.44 46.37 47.48
Mean 12.85 55.72 50.14 54.52
Variance 0.04 0.49 1.75 0.32
@@
993 DMoN–DPR (DVE) 15.43 55.58 47.68 53.96
550 DMoN–DPR (DVE) 14.27 55.01 53.99 58.87
243 DMoN–DPR (DVE) 12.76 56.10 53.71 58.34
16 DMoN–DPR (DVE) 14.62 53.36 49.11 57.45
716 DMoN–DPR (DVE) 13.06 54.46 57.73 62.18
383 DMoN–DPR (DVE) 12.92 55.12 58.39 60.97
277 DMoN–DPR (DVE) 14.04 55.08 51.68 58.26
274 DMoN–DPR (DVE) 12.59 55.68 58.34 60.52
188 DMoN–DPR (DVE) 11.41 57.40 63.84 62.83
796 DMoN–DPR (DVE) 13.51 54.94 48.67 55.26
Mean 13.46 55.27 54.31 58.86
Variance 0.01 0.01 0.27 0.08
Appendix C Code Implementation
The code was implemented by extending the DMoN implementation in PyTorch Geometric (Fey and Lenssen, 2019), and was trained and evaluated using the evaluation protocols found in the official DMoN repository (Tsitsulin et al., 2023).

### v2

@@ E.5 Coauthor Physics
993 DMoN 12.64 59.92 44.49 46.04
550 DMoN 13.87 64.01 54.36 47.54
243 DMoN 13.89 63.77 56.67 49.51
16 DMoN 13.82 63.77 56.89 49.55
716 DMoN 13.93 63.72 55.42 49.02
383 DMoN 13.65 63.63 52.66 46.65
277 DMoN 13.48 63.85 50.84 44.74
274 DMoN 13.84 64.02 53.53 46.36
188 DMoN 13.89 63.73 55.87 49.33
796 DMoN 14.02 64.12 54.29 46.36
Mean $\pm$ Std 13.70 $\pm$ 0.40 63.45 $\pm$ 1.25 53.50 $\pm$ 3.67 47.51 $\pm$ 1.73
@@
993 DMoN–DPR (D) 14.00 62.17 51.59 47.84
550 DMoN–DPR (D) 19.30 56.61 50.82 46.32
243 DMoN–DPR (D) 18.21 57.77 58.61 50.55
16 DMoN–DPR (D) 16.19 54.39 53.19 55.67
716 DMoN–DPR (D) 16.93 55.94 48.09 49.54
383 DMoN–DPR (D) 19.58 55.58 52.71 46.44
277 DMoN–DPR (D) 14.97 55.73 51.49 53.53
274 DMoN–DPR (D) 15.55 56.74 59.80 55.95
188 DMoN–DPR (D) 14.85 55.96 55.95 54.65
796 DMoN–DPR (D) 14.78 58.04 52.68 49.37
Mean $\pm$ Std 16.44 $\pm$ 1.99 56.89 $\pm$ 2.14 53.49 $\pm$ 3.61 50.99 $\pm$ 3.71
@@
993 DMoN–DPR (V) 12.28 58.05 45.73 45.50
550 DMoN–DPR (V) 11.94 57.38 54.27 54.78
243 DMoN–DPR (V) 13.49 60.85 59.11 53.04
16 DMoN–DPR (V) 13.55 58.65 58.50 56.62
716 DMoN–DPR (V) 11.83 58.84 56.84 53.74
383 DMoN–DPR (V) 14.44 57.84 54.21 55.65
277 DMoN–DPR (V) 12.03 58.71 57.09 53.73
274 DMoN–DPR (V) 14.15 57.02 53.90 55.44
188 DMoN–DPR (V) 13.11 57.40 54.62 53.97
796 DMoN–DPR (V) 15.02 61.14 45.01 45.34
Mean $\pm$ Std 13.18 $\pm$ 1.14 58.59 $\pm$ 1.41 53.93 $\pm$ 4.87 52.78 $\pm$ 4.02
@@
993 DMoN–DPR (E) 13.59 59.64 48.13 45.60
550 DMoN–DPR (E) 11.94 57.60 54.02 54.19
243 DMoN–DPR (E) 13.72 61.90 58.72 51.62
16 DMoN–DPR (E) 13.53 58.94 57.15 55.16
716 DMoN–DPR (E) 14.17 57.80 47.71 51.12
383 DMoN–DPR (E) 14.24 58.41 53.68 54.15
277 DMoN–DPR (E) 11.57 58.08 53.16 52.07
274 DMoN–DPR (E) 13.26 57.99 53.77 51.77
188 DMoN–DPR (E) 14.08 60.75 55.54 50.03
796 DMoN–DPR (E) 14.81 62.19 46.39 45.17
Mean $\pm$ Std 13.49 $\pm$ 1.02 59.33 $\pm$ 1.72 52.83 $\pm$ 4.13 51.09 $\pm$ 3.39
@@
993 DMoN–DPR (DV) 15.37 55.95 48.78 53.49
550 DMoN–DPR (DV) 14.84 55.32 52.13 57.57
243 DMoN–DPR (DV) 13.82 56.19 56.71 57.69
16 DMoN–DPR (DV) 11.67 56.14 54.87 57.53
716 DMoN–DPR (DV) 10.97 57.41 62.02 61.81
383 DMoN–DPR (DV) 13.45 56.41 54.45 57.87
277 DMoN–DPR (DV) 14.25 56.26 52.86 57.93
274 DMoN–DPR (DV) 11.79 56.58 57.37 59.70
188 DMoN–DPR (DV) 14.17 56.86 60.24 58.29
796 DMoN–DPR (DV) 13.98 57.70 58.95 58.04
Mean $\pm$ Std 13.43 $\pm$ 1.46 56.48 $\pm$ 0.70 55.84 $\pm$ 4.02 57.99 $\pm$ 2.06
@@
993 DMoN–DPR (DE) 13.94 61.00 47.60 46.37
550 DMoN–DPR (DE) 14.11 56.74 54.67 55.98
243 DMoN–DPR (DE) 13.89 57.00 60.05 58.60
16 DMoN–DPR (DE) 13.90 54.44 53.68 57.90
716 DMoN–DPR (DE) 11.35 57.47 63.18 61.39
383 DMoN–DPR (DE) 15.41 54.93 51.58 53.92
277 DMoN–DPR (DE) 15.84 54.43 51.48 57.57
274 DMoN–DPR (DE) 14.14 57.42 53.66 56.28
188 DMoN–DPR (DE) 15.33 56.20 56.80 55.88
796 DMoN–DPR (DE) 13.83 59.24 47.47 46.26
Mean $\pm$ Std 14.17 $\pm$ 1.24 56.89 $\pm$ 2.09 54.02 $\pm$ 5.00 55.02 $\pm$ 4.99
@@
993 DMoN–DPR (VE) 8.51 36.26 14.21 44.28
550 DMoN–DPR (VE) 13.11 57.01 52.67 57.06
243 DMoN–DPR (VE) 14.40 59.91 56.09 52.22
16 DMoN–DPR (VE) 13.97 58.17 62.31 59.21
716 DMoN–DPR (VE) 13.66 58.24 53.74 53.47
383 DMoN–DPR (VE) 14.12 57.60 56.52 57.86
277 DMoN–DPR (VE) 16.66 60.87 48.23 44.09
274 DMoN–DPR (VE) 12.18 56.56 51.47 56.39
188 DMoN–DPR (VE) 11.16 56.05 56.50 60.71
796 DMoN–DPR (VE) 13.32 58.12 47.75 50.50
Mean $\pm$ Std 13.11 $\pm$ 2.17 55.88 $\pm$ 7.04 49.95 $\pm$ 13.27 53.58 $\pm$ 5.86
@@
993 DMoN–DPR (DVE) 14.25 57.43 42.12 46.26
550 DMoN–DPR (DVE) 14.53 55.66 53.54 57.87
243 DMoN–DPR (DVE) 12.50 55.96 54.67 59.77
16 DMoN–DPR (DVE) 13.27 54.52 51.00 57.72
716 DMoN–DPR (DVE) 12.95 54.76 54.90 58.06
383 DMoN–DPR (DVE) 12.76 55.00 56.93 60.32
277 DMoN–DPR (DVE) 14.47 54.87 52.66 59.00
274 DMoN–DPR (DVE) 12.47 55.39 58.52 61.38
188 DMoN–DPR (DVE) 11.29 57.23 62.59 62.06
796 DMoN–DPR (DVE) 9.90 53.89 48.04 57.18
Mean $\pm$ Std 12.84 $\pm$ 1.45 55.47 $\pm$ 1.14 53.50 $\pm$ 5.67 57.96 $\pm$ 4.42
Appendix F Implementation Details
The code was implemented by extending the DMoN implementation in PyTorch Geometric (Fey and Lenssen, 2019), and was trained and evaluated using the evaluation protocols found in the official DMoN repository (Tsitsulin et al., 2023). The experiments were run on an A100 GPU with 40GB of memory offered by Google Colab Pro. Regarding the runtime analysis results and clustering visualization, the code was run on an Apple M2 Max CPU.